The supplement of an angle is 26 more than five times its complement. what is the measure of your angle?
Answer by oberobic(2304)

A = the angle
C = its complement
C = 90-A
S = its supplement
S = 180-A
.
S = 5*C + 26
.
S = S
so
180-A = 5*C+26
.
substitute C = 90-A
.
180-A = 5(90-A) + 26
.
180-A = 450 -5A + 26
.
180-A = 476 -5A
.
4A = 296
.
A = 74
.
C = 90-74 = 16
.
S = 180-74 = 106
.
Check to see if this solution is the answer.
.
5C + 26 = 5(16) + 26 = 106
Correct.
.
Answer: The angle is 74 degrees.
.
Done.